OpenStudy (anonymous):

solve the equation: log3 X+log3 (x-24)=4 p;ease show me step by step

OpenStudy (anonymous):

base 3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\log _{3}x+\log _{3}(x-24)=4\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x(x-24)=3^4\]can u solve from now?

OpenStudy (anonymous):

oh.. ok.. thanks i see its now:)...

OpenStudy (anonymous):

\[x^2-24x-81=0\]

OpenStudy (anonymous):

(x+3)(x-27)=0

OpenStudy (anonymous):

x=-3 x=27 x=27 is solution

OpenStudy (anonymous):

and i so confuse about.. why dont we take the -3???

OpenStudy (anonymous):

just the formula,look on the lesson loggarithms can't not negative

OpenStudy (anonymous):

ok thanks:)....

OpenStudy (anonymous):

or you can plug x=-3 and x=27 and check